OpenStudy (anonymous):

WIll fan and medal! Solve the following equation: \sqrt{x+4}=3x-12

OpenStudy (anonymous):

x = 5

OpenStudy (anonymous):

but how?

OpenStudy (anonymous):

you could graph it or solve the equation

OpenStudy (anonymous):

i uh, need help with the steps to solve the equation :P

OpenStudy (anonymous):

okay work me through what you have done so far

OpenStudy (anonymous):

\[\left( \sqrt{x+4}^2 \right) = \left( 3x-12 \right)^2\] x+4 = 9x - 144

OpenStudy (anonymous):

your close, you made a mistake simplifying the right hand side of your equation, it should be: (3x-12)(3x-12)

OpenStudy (anonymous):

k, so i got x + 4 = 9x^2 - 72x + 144

OpenStudy (anonymous):

good

OpenStudy (anonymous):

now subtract 4 and x from both sides

OpenStudy (anonymous):

so you get 0 on the left hand side of your equation

OpenStudy (anonymous):

0 = 9x^2 - 72 + 140

OpenStudy (anonymous):

and then i graph that?

OpenStudy (anonymous):

nope you made a mistake

OpenStudy (anonymous):

where did i go wrong..?

OpenStudy (anonymous):

remember you have to foil

OpenStudy (anonymous):

it should be: 0 = 9x^2 - 73x + 140

OpenStudy (anonymous):

you forgot to subtract the x

OpenStudy (anonymous):

then all you have to do is plug it into the quadratic formula

OpenStudy (anonymous):

\[x = \frac{-73\pm \sqrt{(-73)^2 - 4 (9) (140)}\ }{ 18 }\]

OpenStudy (anonymous):

k, that;s how far i got

OpenStudy (anonymous):

yup now solve both the equations with the + and the -

OpenStudy (anonymous):

i keep on getting fractions as answer when i type it in the calc..

OpenStudy (anonymous):

one awnser should be a fraction

OpenStudy (anonymous):

the other should be 5

OpenStudy (anonymous):

i got -5...

OpenStudy (anonymous):

nope its 5

OpenStudy (anonymous):

lol alright :) thanks for your help!